Home, 2024 2023 2022 2021 2020 2019 2018 2017 2016 2015 2014 2012 2011 2010 2009 2008 2007 2006 2005

Haddow Internet Network Status History For 2013

10:30 09-Dec-2013 We have updated the netstatus page, it is now part of the main site, this makes it easy to update it and keep people informed.
14:30 02-Dec-2013 We had a email program glitch on the main mail server, this forced us to block all email access for 20 min, we have fixed the problem and all should be working fine.

15:00 13-Nov-2013 We have had reports from upstream provider that there is some problems in usa with slow links, we are not seeing any problem but some people may have problems, we are watching all links and report if we see a problem.

23:50 31-Oct-2013 Server2 main link back online, the modem on site needed a reboot to bring the line back up.
23:20 31-Oct-2013 Server2 main link back down, we are looking into the problem.
09:00 30-Oct-2013 All Servers, there has been a very heavy load on all servers over the last few days, this has been caused by A LOT of spam hitting the servers, Most of the spam has come from China, the servers are blocking about 98% of all the spam.
07:00 06-Oct-2013 All servers have changed to AEDST, There may be some sites showing the old time due to hard programing of time zones, but these will be changes with in the day.
10:00 01-Oct-2013 Daylight savings starts on the 6th of Oct 2013 at 2am in Aust, the clocks will go forward to 3am and be at AEDST.

02:00 21-Sep-2013 Server2 Main Line Back online, no update as to the problem just there was a problem (That is what we was told) we have checked other companyes and found they had the same problem, But for now the link is up.
19:30 20-Sep-2013 Server2 main line Update, talked to up stream provider who clamed there was no problem, after a few calls they traced the problem as a wide spread problem, No eta on when they might get it fixed.
18:45 20-Sep-2013 Server2 Main link down, Looking into the problem.

14:30 30-Aug-2013 Server2 problems, the links came back up but then went slow, then for some reasion the server went in a panic mode, the server has now been rebooted and all back online, we are looking into why the server has been playing up.
09:45 30-Aug-2013 Server2 link back online, Support needed to attend the site to fix some problems.
22:50 29-Aug-2013 Server2 link down, Both links to server2 are down, we are looking into why.
14:40 28-Aug-2013 Server4 update, there was a DOS attack on one of the core routers all is fixed now.
14:20 28-Aug-2013 Server4 internet slow, looking into the problem.
15:05 15-Aug-2013 Server2 site power, Site back on mains power, they finished ahead of time.
10:15 15-Aug-2013 Server2 site power, Site running on generator, they started a bit late.
10:45 04-Aug-2013 Server1 update, this server is now off, all dns was changed last night, so every one should be on the new server by now.
10:00 01-Aug-2013 Server1 update, the only way to fix the problem is to upgrade, so we have shut the server down.
09:00 01-Aug-2013 Service2 site power, We have been told there will be a power outege on the 15 Aug 2013 from 09:00 to 16:00 for more mains upgrades, the site will run on generator back duing this time.

19:45 31-Jul-2013 Service1 update, there is a glitch in a bit of software on the server & also there was a DDOS attack, which is causing the server to hang, we was looking at shuting down the server in a few weeks but might have to shut it off soon, we are keeping a watch on the server to see if it fixes up, but we are not holding our breath.
19:30 31-Jul-2013 Service1 is not responding, we are looking into the problem.
15:40 14-Jul-2013 Service2 Main Link back online, after a few words and emails to and from, the link is back online, there is no update from the provider about the problem, they still clame the problem was not there, even tho it was.
14:55 14-Jul-2013 Service2 Main Link update, talked to the provider and they clame there is no problem, EVEN tho there is no BGP access to there routers on several ip address.
13:45 14-Jul-2013 Server2 Main link down, talking to upstream provider.
09:00 01-Jul-2013 Services moving, over this month we will be moving all the services that was hosted on Server2 accross to Server4. This new server hosts IPv4 & IPv6 services. there will be no down, but we will be testing sites first before they come live on the new server.

15:45 26-Jun-2013 Server4 installed, dns & web are up all ready, all web sites on server1 will get moved over to this server over the next few weeks, there will be no downtime for websites in the move, We will test the site on this server before it is moved.
15:45 08-Jun-2013 Server1 back online, there was a core cable problem and it took some time to trace down why and how to fix it, backup links were still working fine.
14:45 08-Jun-2013 Server1 link is down, was not meant to but something has happen, hope to have it back online ASAP.
09:00 01-Jun-2013 Server1 core network maintenance, has been reported to be done on the 8 of Jun from 1am UTC time to 4am UTC time, they clame there is not meant to be any down time.

10:00 26-May-2013 server2 server is back online, it took longer to fix due to the HDD failing so we have changed the drive and all back online.
07:37 26-May-2013 server2 server is down, one of the backup HDD ave failed and caused the server to panic and reboot, server is now check all files for damige and hope to be back on line in 30 min, office phone lines are down too bacuse of this.
09:02 24-May-2013 Server2 Main Link, there has been a glitch on this link since 23:50 last night, the link was up but reported as down, had been on the phone to the upstream provider, and it took some time to trace down the problem on the Sydney core router, it was not just us, others comapnys that deal with them had the same problem. In the end the router was given a reboot and that fixed the problem.
15:37 20-May-2013 server2/Office site is back on mains power, 6hr 47min.
08:50 20-May-2013 server2/Office site is running on generator backup.
08:45 20-May-2013 Report from SP Ausnet, that there will be a Mains interruption on server2/Office site from 08:45 to 16:00. WEATHER PERMITTING
Site will be running on generator during this interruption.

08:30 30-Apr-2013 Office phone lines are playing up, we needed to reboot the phone system, it looks like it started the same time as the adsl outege.
07:35 30-Apr-2013 Server2 main link back on line, Link has been down since 1:30am but admin was trying to get some sleep, backup link was working fine..
09:00 07-Apr-2013 All servers. Daylight savings has ended, all servers have reverted back to AEST, some web sites runing php will revert back in a few hrs
00:00 01-Apr-2013 All servers, Daylight savings ends on all servers on the 7 of Apr 2013 at 3am all servers will revirt back to aest (2am).

17:00 31-Mar-2013 Office Phone Lines, back to normal, we have changed the dsl routers firmware and this has the lines back to normal, we are looking at upgreading the dsl router ASAP.
12:00 31-Mar-2013 Office Phone Lines, we have put in a workround for now to get the lines back up and running.
12:00 29-Mar-2013 OFFICE PHONE LINES, we are seeing a problem with all the office lines, calles are going out but nothing coming in, It has been reported to the provider but due to the long weekend it may take time to fix it.
15:00 27-Mar-2013 We have seen some dns problems over the last day or two, it has taken time but we have fixed the problem it only was one domain, all others was working fine.
08:00 20-Mar-2013 Server2 update, server is still fsck the HDD (This server has over 3Tb of HDD), it looks like it just froze up at 3am and to fix the problem we had to do a hard reboot, we are still trying to trace down what had caused the problem.
07:45 20-Mar-2013 Server2 back on line but will be slow for 1hr for HDD checking.
07:30 20-Mar-2013 Server2 is down, we are looking into the problem.
08:45 06-Mar-2013 Server2 links back on line, for some reasion unknown to support & upstream provider, the main link was getting the ipaddress of the backup link, but it was not just H.I. problem, anyone that had 2 or more links with this upstream provider was geting the same problem.
03:45 06-Mar-2013 Server2-backup link down, but both routers showing online, trying to trace the problem.
03:38 06-Mar-2013 Server2 main link down, looking into the problem.
19:00 03-Mar-2013 Due to some problems with the VOIP server, we have had to force the main link to server2 back to the old settings and not via the core router onsite, but we hope to fix this problem with another link router soon.
12:15 02-Mar-2013 All servers, there seem to be an DNS Error that happend on the 28feb2013 on the marster dns server (Server2), this error has been fixed and emails are starting to flow again, some emails may be slow to send or receive for the next hour or so, but will even them self out.
11:45 02-Mar-2013 All servers, there is alot of emails sitting on most of the servers, we are looking into why.

15:15 23-Feb-2013 Server2 links are all up are running on the new routers, we are monitoring all the links just to make sure, Server2 main link working all ok, Backup link1 is having a fit but will fix its self in a few hr's it is online but showing the wrong IP in the DNS.
11:45 23-Feb-2013 Server2, we are starting the upgrade on the routers.
20:00 21-Feb-2013 New updated router equipment has been arriving over this week for the Main office / Server2 links, once configured, we will install it, we hope to do it Saturday the 23rd February, during the upgrade the links may go up and down, or might just not respond, this will be due to some conflicts in the setting up, we will try to have very minimal down time, if there is a major problem we will revert back to the old setup.

14:01 07-Jan-2013 Server2 back on mains, power outge coverd allmost all of Wodonga Shire.
11:45 07-Jan-2013 Server2 site now running on generator.
11:44 07-Jan-2013 Server2 down, no power to site.

All time stated above are in AEST (or AEDST) unless other wise stated. UTC Time Definition, CST Time Definition, AEST Time Definition